OpenStudy (anonymous):

Identify the range of the equation y = −x2 − 8x − 10. All real numbers y ≥ −2 y ≤ 6 y ≤ −4

OpenStudy (anonymous):

since this is a parabola that opens down the range will be from minus infinity up to the second coordinate of the vertex. the first coordinate of the vertex is \(-\frac{b}{2a}\); in your case it is \(-\frac{-8}{2\times -1}=-4\) the second coordinate is what you get when you replace \(x\) by \(-4\)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw